OpenStudy (anonymous):

Write an equation for the line that is parallel to y=3x+8 and passes through the point (-5,1)

OpenStudy (anonymous):

parallel means that the lines have the same slope slope=3 formula y-y=m(x-x) you have a point (-5,1) means (x,y) do you know what to do from here?

OpenStudy (anonymous):

I'm so confused.

OpenStudy (anonymous):

y-1=3(x+5) y-1=3x+15 y=3x+16

OpenStudy (anonymous):

do you get it now and understand what I did?